A block of material of volume 20cm³ and density 24g/cm³ is suspended from a spring balance with half the volume the block immersed in water. What is the reading of the spring balance? (Density of water = 1.0g/cm³)
1
Expert's answer
2019-05-30T09:10:19-0400

The weight of the block immersed in water

"W=mg-g\rho_{\rm{w}} V_{\rm{im}}"

So

"W=g\rho_{\rm{b}}V_{\rm{b}}-g\rho_{\rm{w}}V_{\rm{b}}/2"

"=9.8\:\rm{N/kg}\times 24000\:\rm{kg/m^3}\times 20\times 10^{-6}\rm{m^3}"

"-9.8\:\rm{N/kg}\times 1000\:\rm{kg/m^3}\times 20\times 10^{-6}\rm{m^3}/2"

"=4.6\:\rm{N}"